OpenStudy (anonymous):

Order Of Magnitude Determine whether F(x) grows faster than x^2, same rate as x^2 or slower than x^2 as x approaches infinty. f(x)=x^3+3

OpenStudy (anonymous):

If \(f(x)=x^3+3\) grows faster than \(g(x)=x^2\), then \(f'(x)>g'(x)\), or (for \(g'(x)\not=0\)) \(\dfrac{f'(x)}{g'(x)}>1\). (same rate: change > to =; slower rate: change > to <) Find the limit: \[\lim_{x\to\infty}\frac{f'(x)}{g'(x)}\]

OpenStudy (anonymous):

So I move either x^2 to the denominator and from there I will tell what happens

OpenStudy (anonymous):

suppose \(x=1,000,000=10^6\) then \[(10^6)^3=10^{18}\] whereas \[(10^6)^2=10^{12}\]

OpenStudy (anonymous):

and \[\frac{10^{18}}{10^{12}}=10^6\] so \(x^3\) grows much faster than \(x^2\)

OpenStudy (anonymous):

I get it, so when ever I see a problem like that I can just plug in a big number and solve it.